OpenStudy (anonymous):

solve the equation step by step list each step along with the property you used during that step . -5 x 1 x 11 x 4.

OpenStudy (anonymous):

its just multiplication -220

OpenStudy (anonymous):

thats all it is ?

OpenStudy (anonymous):

ya i mean it was just u were to multiply..........properties i didnt get....... now i guess simply multiply them step by step ... -5*1=-5 -5*11 =-55 -55*4=-220

OpenStudy (anonymous):

ok now i get it thank u very much.